1. The National Institute for Standards and Technology (NIST) mandates that in any store no more than 2% of the items scanned through the electronic checkout scanner should have an inaccurate price. In a recent test at a Wal-Mart a sample of 529 items was randomly selected for scanning. If 2% of the items at this Wal-Mart have an inaccurate price, what is the probability that the sample proportion of items that scan innacurately is between 2.8% and 3.5%?
1
Expert's answer
2019-10-29T11:02:29-0400

Let "X=" the number of successes, "X\sim B(n, p)."

Given that "n=529, p=0.02." Then


"np=529(0.02)=10.58>10"

"n(1-p)=529(1-0.02)=518.42>10"

We can use the normal model


"N(p, \sqrt{{p(1-p) \over n}})=N\big(0.02, \sqrt{{0.02(1-0.02) \over 529}}\big)="

"=N(0.02, 0.0061)"

to approximate the sampling distribution of "\hat{p}".


"P(0.028<\hat{p}<0.035)=P({0.028-0.02 \over 0.0061}<z<{0.035-0.02 \over 0.0061})\approx"

"\approx P(z<2.4590)-P(z<1.3115)\approx"

"\approx0.99303-0.90516\approx0.0879"

The probability that the sample proportion of items that scan innacurately is between 2.8% and 3.5% is "0.0879."